OpenStudy (anonymous):

The radius of a right circular cylinder is increasing at the rate of 3 ft/sec, while the height is decreasing at the rate of 6 ft/sec. At what rate is the volume of the cylinder changing when the radius is 15 ft and the height is 10 ft?

OpenStudy (tkhunny):

Formula for the volume fo a right circular cylinder, please...

OpenStudy (anonymous):

\[v=pir^2h\]

OpenStudy (anonymous):

\[v=\pi r^2h\]

OpenStudy (tkhunny):

That's good. Now, rewrite with everything as a function of time. \(v(t) = \pi \left(r(t)\right)^{2}h(t)\) You don't HAVE to do that. You can just think it. Find the complete derivative, \(\dfrac{dv}{dt}\). This is the calculus part. The rest is algebra.

OpenStudy (dumbcow):

\[\frac{dV}{dt} = \frac{dV}{dr}*\frac{dr}{dt} + \frac{dV}{dh}*\frac{dh}{dt}\] \[ = (2 \pi r h)\frac{dr}{dt} + (\pi r^2) \frac{dh}{dt}\] Sub in all the given conditions for r,h, dr/dt , dh/dt

OpenStudy (tkhunny):

I guess we'll never know if the OP can do it.